OpenStudy (anonymous):

log(base2) x - log(base2) (x-1) = 2

OpenStudy (unklerhaukus):

\[\log_b(A)-\log_bC)=\log_b(A/C)\]

OpenStudy (anonymous):

x/(x-1) = 4 where x is not 1 so x = 4x-4, x = 4/3

OpenStudy (anonymous):

im not quite sure how you knew that x/(x-1) = 4 @LeeYeongKyu

OpenStudy (anonymous):



OpenStudy (anonymous):

oh i get it now thankyou